Circular No. 8784

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006qs–2006sr Fifty-one supernovae discoveries have been reported from survey im-

ages: 2006qs–2006ry by the “Nearby Supernova Collaboration” (details on CBET 768; magnitude range R = 16.1–20.2), 2006rz by the Sloan Digital Sky Survey II collaboration (communicated by J. Frieman; details on CBET 770; magnitude g = 20.0), and 2006sa–2006sq by the ESSENCE project (communicated by P. Challis; details on CBET 772; magnitude range R = 20.8–23.5). All are type-Ia supernovae except 2006qt, 2006rc, and 2006rx (type IIn); 2006rh, 2006rj, 2006rk in MCG +05-17-19, 2006rp, 2006rq, 2006rt, 2006ru, and 2006ry (type II); 2006sa (type IIp); and 2006sg (type Ic). SN 2006rs appears to be associated with IC 28. Those objects brighter than mag 18.8 at discovery are tabulated below. SN 2006sr was discovered by D. Rich (Hampden, ME, U.S.A.) on unfiltered CCD frames taken with a 0.31-m reflector. Additional magnitudes for 2006sr in UGC 14: Oct. 1991, [20.2 (Digitized Sky Survey, red); Aug. 1993, [19.9 (DSS, blue); 2006 Nov. 21.99 UT, [18.6 (Rich); Dec. 15.03, 16.9 (Rich).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006qw Oct. 21.5  3^{h}08^{m}46.^{s}88 +18^{o}56′26′′.0 17.9  –
2006rb Oct. 24.2 23 35 25.09 + 7 06 46.4 17.7  –
2006rk Nov. 6.4  7 16 24.51 +27 37 25.5 16.1  –
2006rp Nov. 9.5  4 09 05.52 +34 38 15.6 18.3  –
2006rs Nov. 11.2  0 33 09.13 −13 27 16.2 16.5  –
2006sr Dec. 12.95 0 03 35.02 +23 11 46.2 17.4 2′′.5 W, 16′′.6 S

COMET P/2006 W1 (GIBBS) Additional observations of this comet (cf. IAUC 8775) have shown it

to be of short period, the elements below taken from MPEC 2006-X40:

T = 2006 Mar. 25.784 TT  ω = 229.^{o}906
e = 0.68718  Ω = 153.814 }2000.0
q = 1.70086 AU  i = 19.101
a = 5.43716 AU n^{o} = 0.077740 P = 12.7 years

COMET 182P/LONEOS The recovery of comet P/2001 WF_{2} = 2006 W2 (cf. IAUC 8776) has

permitted its numbering as comet 182P/LONEOS (cf. MPC 58065).

2006 December 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green